whats the equation of a line that satisfies the given conditions. Through (-9, -7); perpendicular to the line passing through (-6, 5) and (-2, 3)
Answer by Alan3354(69443)

whats the equation of a line that satisfies the given conditions. Through (-9, -7); perpendicular to the line passing through C(-6,5) and D(-2,3)
---------------
Step 1, label the points.
the line passing through A(-6,5) and B(-2,3)
Step 2, find the slope of the line AB
Call it m1
---
Lines perpendicular have a slope that's the negative inverse, = -1/m1
Call it m2.
---
Use y -y1 = m2*(x - x1) where (x1,y1) is a point.
----
The perpendicular line might go thru C and D, might not.